OpenStudy (anonymous):

Two joggers start from different locations and simultaneously begin heading toward each other. One of the joggers jogs 16 mph, while the other jogs 21 mph. If the two joggers are 444 miles apart, how many hours will it take before they meet?

OpenStudy (anonymous):

@Venny

OpenStudy (anonymous):

It is a physics related question. You will have to use relative motion. Velocity of A - Velocity of B = Velocity of AB Frame that is we will keep B at rest and let the other fellow's speed be added up. New Velocity = 16 + 21 = 37 Time = Distance / Speed Time = 444/37

OpenStudy (anonymous):

3.89 ??

OpenStudy (anonymous):

wait nooo

OpenStudy (anonymous):

12 hours

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

I think you could write this as an algebra problem too... 16n+21n=444

OpenStudy (anonymous):

It is easy to grasp in physics.